OpenStudy (anonymous):

assume that 4 molecules of nitrogen and 9 molecules of hydrogen are present? after complete reaction, how many molecules of ammonia are proceded?

OpenStudy (anonymous):

@Lena772

OpenStudy (anonymous):

@AravindG

OpenStudy (anonymous):

@GavinxFiasco

OpenStudy (anonymous):

@marylou004

OpenStudy (anonymous):

@GTXMUQSIT assume that 4 molecules of nitrogen and 9 molecules of hydrogen are present? after complete reaction, how many molecules of ammonia are proceded? can u plz solve this prob

OpenStudy (anonymous):

So \[2N _{2} +3 H _{2} \rightarrow 2 NH _{3}\] that's our balanced equation 4 molecules of N2 would react completely with 4 x 3 = 12 molecules of H2, but there are not that many H2 molecules present, so H2 is the limiting reactant and N2 is in excess. \[9 mol H _{2} \times \left(\begin{matrix}2NH _{3} \div\ 3H2\end{matrix}\right)\] = 6 molecules of ammonia (this can all be seen from out balanced equation)

OpenStudy (gtxmuqsit):

yeah 6 moles